OpenStudy (anonymous):

1. Which equation could be used to solve the question below? Cheryl had 24 charms. She used c charms to make a bracelet. Then, she went to the store and bought 13 more charms. Now, she has 29 charms. How many charms did Cheryl use to make the bracelet? 24 - c - 13 = 29 24 + c - 13 = 29 24 - c + 13 = 29 24 + c + 13 = 29

OpenStudy (anonymous):

Fan and medall

OpenStudy (anonymous):

it simple right

OpenStudy (anonymous):

Cherly had 24 charms Start with 24 Used c charms, so it would subtract -c Bought 13 more charms +13

OpenStudy (anonymous):

so wats the aswer

OpenStudy (anonymous):

You tell me.

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Which equation could be used to solve the question below? Mac picked 26 ears of corn. He sold c at the farmer's market. Then, he sold 13 ears of corn at the produce stand. Now, he has 4 ears of corn left. How many ears of corn did Mac sell at the farmer's market? 26 + c + 13 = 4 26 - c - 13 = 4 26 + c - 13 = 4 26 - c + 13 = 4

OpenStudy (anonymous):

Same like before. Picked 26 ears of corn 26 Sold c -c Then he sold 13 more -13 Have 4 ears left =4

OpenStudy (anonymous):

b

OpenStudy (anonymous):

What about b?